OpenStudy (anonymous):

Theorem: Let a and be be real numbers. If (a+b)^2 <= 6ab then a =b. Proof: suppose that a=b. Then (a+b)^2 =4a^2 and 6ab +6a^2, and since 4a^2 <= 6a^2 is true, the theorem is true. Is the proof correct and is the theorem correct? giving clear reasons for each

OpenStudy (confluxepic):

Welcome to Open Study.

OpenStudy (anonymous):

I think you have to use induction to prove

OpenStudy (anonymous):

Basic step: Suppose it is true for the case a =b =1 \(P(1) := (1+1)^2 \leq 6*1*1\\~~~~~~~~~~~~~~~~~~~~~~~~~~~2\leq 6\) Hypothesis step: Assume it is true for a = b =n That is \[P(_n):= (n+n)^2\leq 6n*n\\~~~~~~~~~~~~~~~~~~~~~~~~4n^2\leq 6n^2\]

OpenStudy (anonymous):

Now, prove \(P_{n+1}\) which is \[P_{n+1}:= ((n+1)+(n+1))^2\leq 6*(n+1)*(n+1)\]

OpenStudy (anonymous):

I don't think you can't handle it. :)

OpenStudy (freckles):

you know you only suppose the if part... if you wanted to try to prove the contrapositive though: you could assume a doesn't equal b and show (a+b)^2>6ab. If this "theorem" is false, then we should be able to find a counterexample. (a+b)^2 will be positive or zero for all real a and b 6ab will be negative for some real values a and b so I think you could actually find a counterexample for (a+b)^2<=6ab

OpenStudy (anonymous):

@OOOPS a proof by induction doesn't work here. We're working with real numbers here. There is no basis case.

OpenStudy (ikram002p):

its if then statement , \(P\rightarrow Q \) there is 3 ways to prove :- 1- direct proof , assume P is true to reach Q is true 2- contrapositive assume not Q in true to reach not P is true 3- contradiction , assume what ever u want to be true but at the end ur result would contradict ur assumption what do you wanna use ? ( im assuming cu i dint find a ounter example xD)

myininaya (myininaya):

I haven't heard anything from @jbn123 are you there?

OpenStudy (anonymous):

This is what i did: Suppose that (a+b)^2 <= 6ab, then a^2+2ab+b^2 <= 6ab and a^2-4ab+b^2<=0 then i wanted to let a=b, so that you eend up with b^2-4b^2+b^2<=0 so -2b^2<=0 but it seems incorrect

OpenStudy (anonymous):

I'm not sure if the theorem is correct or not, i was looking for a counter example. But surely by letting a=b, then the theorem would then become 4a^2 <= 6a^2 which meas for any value of a, the theorem will be true?

OpenStudy (gorv):

yeah its true

OpenStudy (gorv):

because what we considered initially has led to the right result or proved us right by satisfying the condition ...so our assumption and theorm both are true

OpenStudy (anonymous):

The statement is false. Take \(a=-2\) and \(b=2\). Then you have: $$ (-2+2)^2\le 6(-2)(2)\Longrightarrow 0\le -24,$$ which is false.

OpenStudy (gorv):

lolll a=b we have to take

OpenStudy (gorv):

2 is not equal to -2 is 2=-2???

OpenStudy (anonymous):

The statement is, "If \((a+b)^2\le 6ab \), then \(a=b\). You can't assume \(a=b\) to prove the statement.

OpenStudy (anonymous):

You are trying to prove that \((a+b)^2\le 6ab\) forces \(a=b\), which is not the case since I provided a counter example.

OpenStudy (gorv):

they said if (a+b)^2<=6ab

OpenStudy (gorv):

than a=b

OpenStudy (anonymous):

If the problem had been, "Prove if \(a=b\), then \((a+b)^2\le 6ab\).", then you start by assume \(a=b\).

OpenStudy (gorv):

you get 0<=-24

OpenStudy (gorv):

which is not true

OpenStudy (gorv):

means a not equal to be

OpenStudy (gorv):

so the theorm is true

OpenStudy (gorv):

@nerdguy2535

OpenStudy (anonymous):

Let's make sure we agree what the problem wants us to prove. Using what you posed earlier, this is a "\(P\longrightarrow Q\)" problem, with $$P=(a+b)^2\le 6ab$$and $$Q=(a=b).$$ Do we agree on this?

OpenStudy (gorv):

no no no

OpenStudy (gorv):

if P is true than Q

OpenStudy (gorv):

if P comes true than a=b

OpenStudy (anonymous):

Yes. Thats what I typed means. If P is true then Q is also true. Thats the exactly what \(P\longrightarrow Q\) means.

OpenStudy (gorv):

okay....further

OpenStudy (anonymous):

Cool

OpenStudy (anonymous):

Ok, so I claim this statement is false. The statement (as claimed by the original poster) says it should work for all real numbers a and b (The second sentance, "Let a and b be real numbers...). My counter example is letting \(a=-2\) and \(b=2\). For these real numbers, the statement doesn't hold.

OpenStudy (gorv):

check again

OpenStudy (gorv):

there is a word IF before this

OpenStudy (gorv):

which means if P is true

OpenStudy (gorv):

a=-2 and b=2

OpenStudy (gorv):

0<=-24

OpenStudy (anonymous):

AH!

OpenStudy (anonymous):

I understand now. You are definitely right.

OpenStudy (gorv):

thanks @nerdguy2535

OpenStudy (anonymous):

Ok, what about this? \(a=4 ,b=5\)? Then \((5+4)^2=81\) and \(6(4)(5)=120\), so \((a+b)^2\le 6ab\) is true, but, \(4\ne 5\).

OpenStudy (gorv):

applause to u buddy ..u r right now

OpenStudy (gorv):

u proved it wrong

OpenStudy (anonymous):

so rather than try to write a proof for the statement being true or false, we can simply prove it to be false by a counter example?

OpenStudy (gorv):

yeah we can prove it wrong

OpenStudy (gorv):

but by counter example its hard to prove wrong